Initial Setup at Front Bent Gregorian - 18 September 2011 UT

Observer(s): J. Hill (Tucson), O. Kuhn (LBTO) plus Fernando, Mauro, Rick, Rebecca
Telescope Operator: S. Allanson (LBTO)
Telescope Support: M. Midkiff (LBTO)
Instrument Support: T. Shih (LBTO)

Summary

Things go much faster without clouds, rain, humidty! We got on sky at 04:00 after some delays from the power outage.

AzCam was happy, but AGw3 was having probe problems. MCS has a new bug related to rotator unwrapping (and a voice to announce when the bug happens). See IT3643. Temperature was 6 degC during the night with mirrors in decent equilibirum, and telescope 2-3 degrees warmer. Seeing ranged from 0.7 to 1.3 arcsec for most of the night.

We've completed steps 1a-1f from Andrew's list for LFBG. 1e and 1f still have some data analysis by Doug and Andrew needed on Sunday, but the eyeball analysis looks OK. We've not made a new pointing model - this would be useful, but we can live with the old one for now. A new pointing model should be made after the final binodal adjustments are done (or even after range-balanced collimation models).

Take Away: LFBG is back like it was before - no range-balanced pointing or collimation improvements have been implemented.

We are handing over the telescope to Olga for LBC verification for the last couple hours of the night.

Stanford 37 - Arizona 10 (but our TV is bigger!)

Preparations

We did another all day power outage. Things (infrastructure) were mostly back up and running (which is not to say ready for observing) about 02:15 UT.

The monsoon weather is at long last departed! Skies are clear!

Vent doors are open midway and mirror ventilation running as we: fill PISCES, adjust M2 SX swing arm, turn on AO secondaries, fill MODS1, work on LBC flushing, etc.

Michael and Tim adjusted the SX M2 swing arm back to its previous position (the historical correct position).

Details

Open

03:30 Prepare to open shutter doors

03:42 Doors are open.

Initialize M2 hexapods and M3s, set M2 adaptive shells.

IRTC @ RFBG

03:59 Preset to M5 star ACQUIRE - adjust pointing IE=29 CA=6

IRTC @ LFBG

Sun Sep 18 04:07:25.896 2011  TCS authorize alert for left IRTC at bentGregorianFront focus and right NONE at none focus: PresetFlat[8] failed - File not found: IRTC

Using the IRTC_RFBG_20101226 pointing model to start with (what we just used to find the star on the right side). John manually loaded it on the PCS pointing details GUI.

04:08 Preset to M5 (20:38:31 +21:12:04) star ACQUIRE - ptspiral
We found a star, and adjusted focus. But it wasn't the M5 star (too faint). The coma is pleasantly modest.
Why is the star drifting away at ~1 arcsec per second? The same pointing model was OK on the right side.

LUCI1 @ LFBG

04:27 Preset to M5 (20:38:31 +21:12:04) star ACQUIRE - found multiple star images, but not the M5 star. Now the drift is slower. ??

04:31 Preset to other M5 star

04:32 Preset to Vega - diffraction spikes say that we are close, but it is still challenging to find the actual star. Found it at approx IE=-15 CA=165

04:45 preset to M5 star - precise centering is IE=-28 CA=+159 and we need to translate M2 so that IE=-72 CA=64

Previous Values of Global Offsets: Xglob=+0.20 Yglob=+1.00 Z=?? RX=-104 RY=+21
from our translation DX M2 Xglob=4.50 Yglob=3.10 Zglob=+3.00 still with some coma

04:57 preset to collimation star ACTIVE (split mode) - seeing is 0.75 arcsec

DX M2 Xact=-4.248 Yact=-1.054 Zact=-0.060 RX=112.5 RY=-431.1 (for EL=62.5) Sum X=+0.252 Y=+2.046 Z=+3.00 RX=8.5 RY=-410.1

05:05 star at lower elevation - missed the acq field

rotator faulted..............

IE=-36 CA=+52 at EL=36 (without obvious coma or astigmatism with shell RIP'd)

05:20 installed the new collimation model DXSMLUCIFERCollimation.dat

Both shells RIP'd by restarting MCSPU for the rotator issue.
Stray light on WFS sensor was moonlight shining in the vent door.

Correction Data for the SXSMLUCI collimation model

EL=37.6 Xact=0.187 Yact=0.454 Zact=-0.069 RX=-56.6 RY=+26.3
EL=48.7 Xact=0.459 Yact=0.348 Zact=-0.048 RX=-36.3 RY=+35.6
EL=58.7 Xact=+0.513 Yact=-0.237 Zact=-0.050 RX=+24.7 RY=+49.2

Sun Sep 18 05:45:03.759 2011  LFBG rotator tracking has stopped because it was going to hit the limit in 0.000000000000000e+00 sec
But rotator stopped at 335 deg? Steve manually unwraps to -24 deg. Rotator went to 112.8 when Steve resent the preset.

EL=74.4 Xact=+0.515 Yact=-0.734 Zact=-0.072 RX=90.7 RY=63.2
EL=85.8 Xact=0.387 Yact=-1.248 Zact=-0.159 RX=136.7 RY=38.3

IE=-53 CA=+56 at EL=85.5

Field Data on Stone_O

06:06 ACQUIRE preset to Stone_O IE=-53 CA=63 at EL=55

06:15 starting to take field data with IDL field_collect (lots of astigmatism at the initial star)

      preset, Stone_O, MODE=ACTIVE, GS=0, PA=-60.6532
  
% Compiled module: IRC_PRESETTELESCOPE.
% Compiled module: COLLIMATE_GCS.
% Compiled module: COLLIMATE_STATUS__DEFINE.
% Compiled module: WFSC_STRUCT__DEFINE.
% Compiled module: ZERNIKE_VECTORS__DEFINE.
% Compiled module: WFSC_PARAMS__DEFINE.
% Compiled module: IRC_GETPARAMETER.
% Compiled module: IRC_READ_DD_FILE.
% Compiled module: IRC_DD_VARIABLES__DEFINE.
===> Error in irc_getparameter
% Compiled module: IRC_FIND_PARAM.
===> Active Optics loop is not running

Unrestrict DD variables on IIF GUI, then rerun field_collect. That solves the problem for now. See Issue 3644.

           242 nm rms error in image left_wfscimage000050.fits
  => Running collimate_process: /Repository/GCSfiles/dailyimages//left_wfscimage000050.fits
  => Collimated in 2 iterations
  
      # Target star: 0   Stone_O
        21:47:54.0276 +00:00:01.892 RADEC 2000.0 J2000 0.0 0.0 11.63 R 0.00 B_V 1.65
      # Guide star: 21
        21:47:40.4130 +00:01:56.714 RADEC 2000.0 J2000 0.0 0.0 10.82 R 0.00 B_V 0.63
  
      preset, Stone_O, MODE=ACTIVE, GS=21, PA=-60.6532
  
===> Error in irc_presetTelescope
        err = Resmsg[0]  = PresetTelescope result status: Warning         err = Resmsg[1]  = Stop[39] complete
        err = Resmsg[2]  = WARNING: guide cam exposure time is less than 2000 ms which might cause guiding on atmospheric dispersions
        err = Resmsg[3]  = adjusted exposure time to 500 ms
% Program caused arithmetic error: Floating illegal operand

Remove the bright star 21 from the list

06:28 restart field_collect

Now the on-axis star is too bright - seeing is too good or sky is too clear !!!
===> Error in irc_presetTelescope
        err = Resmsg[0]  = PresetTelescope result status: Warning         
        err = Resmsg[1]  = Stop[40] complete
        err = Resmsg[2]  = WARNING: guide cam exposure time is less than 2000 ms which might cause guiding on atmospheric dispersions
        err = Resmsg[3]  = adjusted exposure time to 500 ms

06:30 ACQUIRE preset to BS9107 IE=-36 CA=+64 at EL=36 (yes, it is far over, but the selection of pre-planned fields is quite limited) Seeing is 1.50 arcsec down here, but improving as we rise.

Rotator goofy limit again, Steve unwraps manually. See Issue 3643. The problem looks to be that MCSPU is not recognizing cases where it must unwrap.

Talked to Doug. IDL doesn't aapear to have the correct pupil position parameters, but that can be corrected in post-processing.

      # Target star: 0   BS9107
        03:32:02.6035 +37:20:39.204 RADEC 2000.0 J2000 0.0 0.0 13.56 R 0.64 B_V 1.65
      # Guide star: 28
        03:31:48.5652 +37:23:37.333 RADEC 2000.0 J2000 0.0 0.0 11.22 R 1.23 B_V 0.63
  
      preset, BS9107, MODE=ACTIVE, GS=28, PA=-43.1957
  
===> Error in irc_presetTelescope
        err = Resmsg[0]  = PresetTelescope result status: Warning         err = Resmsg[1]  = Stop[45] complete
        err = Resmsg[2]  = WARNING: guide cam exposure time is less than 2000 ms which might cause guiding on atmospheric dispersions
        err = Resmsg[3]  = adjusted exposure time to 500 ms
% Program caused arithmetic error: Floating illegal operand

07:00 move back to Stone_O while Doug has suppressed IDL from exiting on preset warnings (and maybe also real errors) IE=-51 CA=70 at EL=49
empirically IIF returns 1=error, 2=warning, 0=success

07:04 field_collect

GStar wfsc #
0 102
30 103-105
0 106-110

Now a real error:
  => Collimated in 5 iterations
  
      # Target star: 0   Stone_O
        21:47:54.0276 +00:00:01.892 RADEC 2000.0 J2000 0.0 0.0 11.63 R 0.00 B_V 1.65
      # Guide star: 18
        21:47:39.9517 +00:01:00.476 RADEC 2000.0 J2000 0.0 0.0 12.69 R 0.00 B_V 0.63
  
      preset, Stone_O, MODE=ACTIVE, GS=18, PA=-74.4923
  
result from IIF 1
    setting reval val to 1
===> Active Optics loop is not running
% Program caused arithmetic error: Floating illegal operand

07:18 field_collect, START=3 - the same error moving the probe even to the on-axis position.

Andrew arrives fresh from Munich.

07:20 Steve does stopAGW/startAGW and restarts GCSL

07:30 field_collect, START=2 Seeing fluctuates from 0.8 to 1.3 arcsec at EL=42.

GStar wfsc #
0 111
18 112-114
0 115-116
32 117-119
0 120
20 121-123
0 124
22 125-127
0 128-129
29 130-132
0 133-135

07:57 field_collect on BS9107

Sun Sep 18 08:05:13.147 2011  Preset alert for left LUCIFER mode ACTIVE: guidestar list successfully changed but probe couldn''t be moved to position.

Steve restarts AGw3 again!

Anfrew goes home to get some sleep after 28 hours of travelling.

John restarted IDL here to get Doug's new pupil parameters (updated to match GCS). The on-axis pupil is center at X=8 Y=6, so something is still off somewhere. See wfscimage 0146.

08:13 field_collect on BS9107 at EL=56

GStar wfsc #
0 72-78
23 83-85
0 86-87
GStar wfsc #
0 140-144
GStar wfsc #
0 145-146
23 147-149
0 150-151
28 152-154
0 155
36 156-158
0 159-161
33 162-164
0 165
8 166-168
0 169-171
19 173-175
0 176
42 177-179
0 180
39 181-183
0 184

LBC Verification

08:55 Move to zenith and reconfigure for LBC work (Olga is working in the Science Prep logs)

09:16 Reconfigured and authorized for LBC.

Olga's notes are located in SciPrep20110918

Close

-- JohnHill - 18 Sep 2011
Topic revision: r3 - 20 Sep 2011, JohnHill
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback